5/11 multiplied by 10/9
Found 2 solutions by kekegeter , Alan3354:
Answer by kekegeter (97)

5^2*2/11*3^2
5^1+1*2/11*3^2
50/99

Answer by Alan3354(69443)

5/11 multiplied by 10/9
-------------
(5/11)*(10/9) = (5*10)/(11*9)
= 50/99